OpenStudy (anonymous):

Write each logarithmic equation as a single logarithm (2logbx/3+3logby/4)-5logbz

OpenStudy (anonymous):

please put some brackets for simplification. this seems very obsolete

OpenStudy (anonymous):

(2/4(log(b)x) + 3/4(log(b)y)- 5log(b)z = (log(b)x^(2/4) +log(b) y^(3/4)) -log(b)z^5

OpenStudy (anonymous):

can you explain how to do it

OpenStudy (anonymous):

\[\log_{b}((x^{1/2} * y^{3/4})/z^5) \] this is the correct way.

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

okay. i'll show you

OpenStudy (anonymous):

first of all, what you gave me: \[(1/2(\log_{b} x) +3/4(\log_{b} y) ) - 5\log_{b}z \] correct me if m wrong

OpenStudy (anonymous):

2/3(logbx)

OpenStudy (anonymous):

(2/3) ! stupid of me overlooking that in that case here is how it should look \[((2/3) \log_{b}x + (3/4) \log_{b}y)-5\log_{b}z \]

OpenStudy (anonymous):

which is equal to : \[(\log_bx^{2/3} + \log_by^{3/4})-\log_bz^5\]

OpenStudy (anonymous):

& since their base is common, u can group them like this using the basic rules of logarithm : \[\log_b((x^{2/3} * y^{3/4})/z^5)\]

OpenStudy (anonymous):

\[(\frac{ 2\log_{b}x }{ 3}+\frac{ 3\log_{b} y }{ 4})-5\log_{b} z\]

OpenStudy (anonymous):

yes, i have given the evaluated answer to this equation :)

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

first of all remember that the (2/3) that is multiplied by \[\log_{b} x\] can also be written as \[\log_{b} x^{2/3}\] ie: we are just going in the opposite way of how logarithms are usually provided. for example: \[\log_ba^2 = 2\log_ba\]

OpenStudy (anonymous):

u know this rule right? it is possible to take the power of 'a' in the equation that i mentioned above & multiply it to the log.